OpenStudy (anonymous):

Write the equation of the tangent line to the curve y = ln (2/e^2x) at the point (0, ln2).

OpenStudy (anonymous):

it is really \[\ln(\frac{2}{e^2})\]??

OpenStudy (anonymous):

heh.

OpenStudy (anonymous):

it's a line... so I doubt it.

OpenStudy (anonymous):

\[\ln(\frac{2}{e^{2x}})\] right?

OpenStudy (anonymous):

it is a line

OpenStudy (anonymous):

\[\ln(\frac{2}{e^{2x}})=\ln(2)-\ln(e^{2x})=\ln(2)-2x\]

OpenStudy (anonymous):

line with slope \(-2\) and \(y\) intercept \(\ln(2)\) so whatever you do , do not use any calculus

OpenStudy (anonymous):

This!